# Utility Maximization in Wireless Backhaul Networks with Service Guarantees

### v1

@@ I Introduction
There is a considerable body of work on network utility maximization. The problem was first formulated by Kelly in [kelly1998rate], who proposed allocating rates to users to maximize utility in the network. A distributed solution to this problem was developed in [low1999optimization]. It has been studied extensively in the context of stochastic network optimization, using Lyapunov-based approaches to maximize utility while ensuring network stability [neely2010stochastic]. A summary of cross-layer techniques for wireless networks can be found in [georgiadis2006resource].
@@
There has also been work on utility maximization with service guarantees. Hou and Kumar introduced a framework for meeting service requirements in single-hop wireless networks [hou_theory_2009] and showed that this can be used to maximize utility in [hou2010utility], where utility is determined by the fraction of user traffic which is served before its deadline. This framework was extended to multi-hop networks in [li2012scheduling].
@@
Additionally, there is a growing body of work on scheduling in wireless backhaul networks. In [kabbani_distributed_2007], a throughput-maximizing scheduling policy is developed for wireless backhaul, and in [gopalam2022distributed] a backpressure based policy is proposed with the same objective. In [narlikar2010designing], a scheduling policy is designed which achieves at most twice the delay of an equivalent wired backhaul by multiplexing over both time and frequency. An online learning algorithm is used in [han2018online] to minimize average wireless backhaul delay, while a cross-layer routing, scheduling, and resource allocation design is explored in [cao2007multi].
@@
Offline scheduling for throughput maximization in general wireless networks has also received considerable attention [jain2003impact, hajek_link_1988, kodialam2003characterizing]. When deadline constraints are considered, optimal schedules can be found by relaxing interference and capacity constraints [singh2018throughput, singh2021adaptive]. Without these relaxations, the authors of [djukic2007quality, djukic_delay_2009] show that packet delays can be bounded as a function of the schedule length by solving a mixed-integer program or using heuristic approaches.
@@
Recently, a scheduling framework was proposed to make hard throughput and delay guarantees in wireless networks with interference constraints and general topologies [jones_optimal_2024]. Using a network slicing model, the authors show that tight deadline guarantees can be made by bounding inter-scheduling times for each link on a packet’s route. Constructing schedules with bounded inter-scheduling times is known as pinwheel scheduling and was first introduced in [holte1989pinwheel]. It is known to be NP-complete [mok_algorithms_1989], but polynomial time algorithms exist which can form schedules under certain sets of conditions [chan_general_1992, chan_schedulers_1993, han_distance_1996, lin_pinwheel_1997].
@@
In this work, we build on the network slicing model and inter-scheduling conditions of [jones_optimal_2024] to make simultaneous throughput and delay guarantees for flows in a wireless backhaul network. We introduce a novel state of the art algorithm called Inductive Scheduling ($IS$) that significantly expands the set of pinwheel schedules that can be found in polynomial time. We derive feasibility conditions for pinwheel schedules based on this algorithm, and optimize over these scheduling conditions and flow admission decisions to efficiently maximize utility for a batch of flows, while guaranteeing that each flow’s SLA is satisfied.
@@ II-A Preliminary Results
In [jones_optimal_2024], the authors show that for any wireless topology and interference constraints, and using the same network slicing model described above, deadline guarantees can be made by satisfying conditions on maximum inter-scheduling times $\boldsymbol{k}$. Specifically, they show the following result for a given set of flows $\mathcal{F}$.
Lemma 1 ([jones_optimal_2024]).
@@
This result shows that we can restrict ourselves to cyclic policies without loss of optimality, and shows how to construct an SLA-satisfying policy in terms of slice widths and inter-scheduling times. The deadline constraint (3) combined with the slice width condition (2) guarantees that all deadlines are met by forcing links to be scheduled somewhat regularly and queue sizes to remain small, while the slice capacity condition (4) ensures that a link has enough capacity to support the allocated slices. We encourage readers to reference [jones_optimal_2024] for more details.
Define the class of cyclic, work-conserving scheduling policies that satisfy our local interference constraints as $\Pi_{c}$. In order for queues to remain stable and for any finite deadline to be met as the SLA duration $T$ goes to infinity, the service rate of each queue must be at least as large as the arrival rate [jones_optimal_2024]. Defining the time-average activation rate of link $e$ under policy $\pi$ as
@@ IV Inductive Scheduling
As discussed in the introduction, forming a schedule with constraints on maximum inter-scheduling times is known as pinwheel scheduling, and determining whether such a schedule exists for a given $\boldsymbol{k}$ is NP-complete [holte1989pinwheel]. Polynomial-time algorithms exist which can schedule certain classes of vectors, and the state of the art algorithm, known as $S_{xy}$ in the literature, is able to schedule all known polynomial-time schedulable vectors [chan_general_1992].
@@
We further define the set of all vectors with densities less than or equal to $\rho$ as $\mathcal{K}_{\rho}$. It is easy to see that no vectors outside of $\mathcal{K}_{1}$ are schedulable, which provides an upper bound and a necessary condition on schedulable density. While this condition is far from sufficient, some vectors with density $1$ are schedulable. Consider a vector of $M$ elements, each equal to $M$. Vectors of this form have a density of $1$ and are trivially schedulable by a round-robin policy. More generally, however, sufficient conditions are difficult to obtain. It was recently shown that all vectors in $\mathcal{K}_{5/6}$ are schedulable [kawamura_proof_2024], and this is the largest possible density bound. This can be seen by considering the vector $\boldsymbol{k}=(2,3,x)$, which has density $5/6+1/x$ and is not schedulable for any value of $x$.
@@
then $\boldsymbol{k^{\prime}}$ is schedulable, and as a result $\boldsymbol{k}$ is schedulable. The $S_{xy}$ algorithm, which has complexity $O(M^{2})$, searches for values of $\boldsymbol{k^{\prime}}$, $x$, and $y$ which satisfy this condition, and if found produces a valid schedule. We encourage readers to reference [chan_general_1992] for complete details.
@@ VI-B DSUM Algorithm
We tested the performance of the $DSUM$ algorithm on various trees of depth $3$, while varying the link capacities, the number of children (i.e., the degree) at each node, and the number of flows at each AP. For the purposes of these simulations, we set $\lambda=10$ Mbps for each flow, and the duration of a time slot as $125\ \mu s$, the smallest time slot used for data in the 5G standard [3gpp-ts-138211]. Note that both $\lambda$ and the time slot duration can be scaled without changing the results, provided the link capacities are scaled with $\lambda$.
@@ VII Conclusion
For compactness, we will use vector and matrix notation in this section, denoting vectors in bold and matrices with capital letters. In addition, we make use of the notation $\mathbf{1_{M}}$ where possible to denote a vector of ones with length $M$. Before introducing the $IS$ constraints, we first address the nononvex slice capacity constraint (22d), which has a bilinear quadratic term. To linearize this constraint, we employ the well-known McCormick envelope technique [mccormick1976computability], which introduces additional variables and constraints to linearize bilinear terms when at least one of the variables is binary. We will employ this technique several times going forward. Let $\boldsymbol{\kappa_{e}}$ be the binary representation of $k_{e}$, such that

### v2

@@ I Introduction
There is a considerable body of work on network utility maximization. The problem was first formulated by Kelly in [19], who proposed allocating rates to users to maximize utility in the network. A distributed solution to this problem was developed in [23]. It has been studied extensively in the context of stochastic network optimization, using Lyapunov-based approaches to maximize utility while ensuring network stability [27]. A summary of cross-layer techniques for wireless networks can be found in [7].
@@
There has also been work on utility maximization with service guarantees. Hou and Kumar introduced a framework for meeting service requirements in single-hop wireless networks [14] and showed that this can be used to maximize utility in [13], where utility is determined by the fraction of user traffic which is served before its deadline. This framework was extended to multi-hop networks in [21].
@@
Additionally, there is a growing body of work on scheduling in wireless backhaul networks. In [17], a throughput-maximizing scheduling policy is developed for wireless backhaul, and in [8] a backpressure based policy is proposed with the same objective. In [26], a scheduling policy is designed which achieves at most twice the delay of an equivalent wired backhaul by multiplexing over both time and frequency. An online learning algorithm is used in [11] to minimize average wireless backhaul delay, while a cross-layer routing, scheduling, and resource allocation design is explored in [2].
@@
Offline scheduling for throughput maximization in general wireless networks has also received considerable attention [15, 9, 20]. When deadline constraints are considered, optimal schedules can be found by relaxing interference and capacity constraints [28, 29]. Without these relaxations, the authors of [6, 5] show that packet delays can be bounded as a function of the schedule length by solving a mixed-integer program or using heuristic approaches.
@@
Recently, a scheduling framework was proposed to make hard throughput and delay guarantees in wireless networks with interference constraints and general topologies [16]. Using a network slicing model, the authors show that tight deadline guarantees can be made by bounding inter-scheduling times for each link on a packet’s route. Constructing schedules with bounded inter-scheduling times is known as pinwheel scheduling and was first introduced in [12]. It is known to be NP-complete [25], but polynomial time algorithms exist which can form schedules under certain sets of conditions [4, 3, 10, 22].
@@
In this work, we build on the network slicing model and inter-scheduling conditions of [16] to make simultaneous throughput and delay guarantees for flows in a wireless backhaul network. We introduce a novel state of the art algorithm called Inductive Scheduling ($IS$) that significantly expands the set of pinwheel schedules that can be found in polynomial time. We derive feasibility conditions for pinwheel schedules based on this algorithm, and optimize over these scheduling conditions and flow admission decisions to efficiently maximize utility for a batch of flows, while guaranteeing that each flow’s SLA is satisfied.
@@ II-A Preliminary Results
In [16], the authors show that for any wireless topology and interference constraints, and using the same network slicing model described above, deadline guarantees can be made by satisfying conditions on maximum inter-scheduling times $\boldsymbol{k}$. Specifically, they show the following result for a given set of flows $\mathcal{F}$.
Lemma 1 ([16]).
@@
This result shows that we can restrict ourselves to cyclic policies without loss of optimality, and shows how to construct an SLA-satisfying policy in terms of slice widths and inter-scheduling times. The deadline constraint (3) combined with the slice width condition (2) guarantees that all deadlines are met by forcing links to be scheduled somewhat regularly and queue sizes to remain small, while the slice capacity condition (4) ensures that a link has enough capacity to support the allocated slices. We encourage readers to reference [16] for more details.
Define the class of cyclic, work-conserving scheduling policies that satisfy our local interference constraints as $\Pi_{c}$. In order for queues to remain stable and for any finite deadline to be met as the SLA duration $T$ goes to infinity, the service rate of each queue must be at least as large as the arrival rate [16]. Defining the time-average activation rate of link $e$ under policy $\pi$ as
@@ IV Inductive Scheduling
As discussed in the introduction, forming a schedule with constraints on maximum inter-scheduling times is known as pinwheel scheduling, and determining whether such a schedule exists for a given $\boldsymbol{k}$ is NP-complete [12]. Polynomial-time algorithms exist which can schedule certain classes of vectors, and the state of the art algorithm, known as $S_{xy}$ in the literature, is able to schedule all known polynomial-time schedulable vectors [4].
@@
We further define the set of all vectors with densities less than or equal to $\rho$ as $\mathcal{K}_{\rho}$. It is easy to see that no vectors outside of $\mathcal{K}_{1}$ are schedulable, which provides an upper bound and a necessary condition on schedulable density. While this condition is far from sufficient, some vectors with density $1$ are schedulable. Consider a vector of $M$ elements, each equal to $M$. Vectors of this form have a density of $1$ and are trivially schedulable by a round-robin policy. More generally, however, sufficient conditions are difficult to obtain. It was recently shown that all vectors in $\mathcal{K}_{5/6}$ are schedulable [18], and this is the largest possible density bound. This can be seen by considering the vector $\boldsymbol{k}=(2,3,x)$, which has density $5/6+1/x$ and is not schedulable for any value of $x$.
@@
then $\boldsymbol{k^{\prime}}$ is schedulable, and as a result $\boldsymbol{k}$ is schedulable. The $S_{xy}$ algorithm, which has complexity $O(M^{2})$, searches for values of $\boldsymbol{k^{\prime}}$, $x$, and $y$ which satisfy this condition, and if found produces a valid schedule. We encourage readers to reference [4] for complete details.
@@ VI-B DSUM Algorithm
We tested the performance of the $DSUM$ algorithm on various trees of depth $3$, while varying the link capacities, the number of children (i.e., the degree) at each node, and the number of flows at each AP. For the purposes of these simulations, we set $\lambda=10$ Mbps for each flow, and the duration of a time slot as $125\ \mu s$, the smallest time slot used for data in the 5G standard [1]. Note that both $\lambda$ and the time slot duration can be scaled without changing the results, provided the link capacities are scaled with $\lambda$.
@@ VII Conclusion
For compactness, we will use vector and matrix notation in this section, denoting vectors in bold and matrices with capital letters. In addition, we make use of the notation $\mathbf{1_{M}}$ where possible to denote a vector of ones with length $M$. Before introducing the $IS$ constraints, we first address the nononvex slice capacity constraint (22d), which has a bilinear quadratic term. To linearize this constraint, we employ the well-known McCormick envelope technique [24], which introduces additional variables and constraints to linearize bilinear terms when at least one of the variables is binary. We will employ this technique several times going forward. Let $\boldsymbol{\kappa_{e}}$ be the binary representation of $k_{e}$, such that